# Modeling Time-Variant Responses of Optical Compressors with Selective State Space Models

### v2

@@ 4.3.2 Listening Tests
We use Mushra tests (Series,, 2014) to evaluate the perceptual similarity of the compressed signal produced by different models against the actual reference recording taken from the analog optical compressors. Listening tests are only carried out for models trained on the TubeTech CL 1B and Teletronix LA-2A datasets. The tests are conducted using the WebMUSHRA platform  (Schoeffler et al.,, 2018) and involve only participants with music performance or production expertise. For the test, we utilize a standardized setup consisting of a laptop, studio-grade audio interface, and high-end studio headphones.
Initially, the participants are briefed on the test’s objectives and given instructions on using the listening test’s graphical user interface. Each session comprised six different listening tests conducted in a fixed order. Participants are asked to rate the following eight signals on a scale from 00 to $100$ in comparison to the reference recording: the original ’hidden’ reference signal, the reference signal with low-pass filtering at $7$ kHz and $3.5$ kHz, the outputs of the proposed model with the S6 block, alternative models with recurrent blocks based on S4D, ED, and LSTM, and the output from a commercial software emulation of the specific compressor. The test signals, lasting 5 seconds each, covered a variety of drums, guitar, and piano sounds. Baseline models are not included in the test. Participants are allowed to listen to the audio samples repeatedly to ensure accurate evaluation. Of the six tests, three pertained to models of the Teletronix LA-2A compressor and three to the TubeTech CL 1B.
@@ 5 RESULTS AND ANALYSIS
Table 4 reports the results for models trained on the Teletronix LA-2A and TubeTech CL 1B datasets, detailing loss and quantitative evaluation metrics based on test set performance. The table clearly shows that the proposed model with S6 layers outperforms all others in replicating analog optical compressors’ time-variant nonlinear sound alteration processes—processes that vary according to each device’s specific control parameters. This is particularly true for the TubeTech CL 1B, where the performance gap between the S6 and all the other models is notably larger. In contrast, for the Teletronix LA-2A, the S6 model performs more closely to the S4D, ED, and LSTM models. In both cases, the proposed S6 model achieves lower errors for all considered metrics, with the exception of PSM, which is less indicative due to values close to $1$ for all the models. This suggests minimal differences among the models according to the PEMO-Q auditory perception model and underscores the need for actual listening tests. While S4D and ED models exhibit similar behavior on this modeling task, the LSTM models produced the worst results. A similar trend is observed with LSTM-b among the baselines. Lastly, ED-b demonstrates notably high accuracy for the Teletronix LA-2A dataset.
@@
For the MUSHRA listening test, a total of $20$ individuals participated, with no participant meeting the criteria for result exclusion. The results are illustrated in Figure 6, where the box plots display the median, lower and upper quartiles, maximum and minimum values, and outliers of the responses. Results indicate that S6 models are slightly more perceptual accurate than others, offering the closest match to the reference. This is demonstrated by the median level and the narrower interquartile range, suggesting consistent performance across listening tests. The S4D models appear to be the second best, while the LSTM models performed particularly poorly when modeling the TubeTech CL 1B, likely due to noticeable audible artifacts. The ratings of the equivalent compressed signals generated by the software emulations Universal Audio LA-2A and Softube CL 1B, which are slightly lower than those of most models, should be interpreted with care. These signals, which may sound clearer and thus less similar to the reference, are produced digitally without passing through analog-to-digital conversions and analog circuits, which can introduce a small amount of noise that may be perceptible.
@@
For further qualitative evaluation, we have computed a collection of plots to compare the response of the Teletronix LA-2A and TubeTech CL 1B, used as targets for the trainings, with that of the proposed S6 model and alternative models with recurrent blocks based on S4D, ED, and LSTM. These plots were generated for a selection of compression parameter settings representing cases commonly used in real application scenarios. Due to space limitations, this paper only includes those related to the following settings: Teletronix LA-2A in limit mode with peak reduction set to $20$ and compression mode with peak reduction set to $40$; TubeTech CL 1B with a threshold $-20$ dBu and ratio $2:1$, and with threshold $-10$ dBu and ratio $6:1$, both with attack and release set at $\sim 250$ ms and $\sim 5$ s. The plots we computed include the spectrograms of the output signals in Figure 7 and Figure 8, the RMS of the output providing an indication of the signal level over time in Figure 9, and the spectral flux of the output highlighting the response during sharp attacks in Figure 10. Spectrograms are calculated using windows of $1024$ samples with $25\%$ overlap. The RMS and spectral flux are computed using windows of $4096$ samples with $25\%$ overlap.
Figure 7: Spectrograms of the output of the Teletronix LA-2A and of the output of models, trained using the proposed S6 and alternative architectures. Each column presents spectrograms for different compression settings labeled above the columns. The model that produced each spectrogram is specified at the top of the respective spectrogram. ’Target’ refers to the analog hardware device.
Figure 8: Spectrograms of the output of the TubeTech CL 1B, and of the output of models trained using the proposed S6 and alternative architectures. Each column presents spectrograms for different compression settings labeled above the columns. The model that produced each spectrogram is specified at the top of the respective spectrogram. ’Target’ refers to the analog hardware device.
@@
Figure 9: RMS levels measured at the output of the Teletronix LA-2A (top) and TubeTech CL 1B (bottom) compared with those at the output of models trained using the proposed S6 and alternative architectures. The applied compression settings are indicated above the plots. The legend associates the different color curves with their corresponding models, with ’target’ denoting the analog hardware device.
In replicating the appropriate gain reduction required to match the target compressed signal, the LSTM also shows the poorest performance, as evident in the RMSE plot examples of Figure 9, which refers to a $3$-second snippet from the test set. The snippet features two plucked electric bass notes in the first two seconds, followed by vocals in the final second. The first bass note is muted, while the second, occurring in the middle of the snippet, has a longer sustain. These examples have been selected as they clearly illustrate the differing behaviors of various models compared to the target, especially during the sustained bass note, which presents the greatest challenge, representing a region of the test set with errors notably higher than the average. The concluding vocal segment displays a series of rapid, albeit less pronounced, level changes that several models fail to compress accurately. It appears that most models follow the amplitude envelope of the output signal, yet they vary in their degree of accuracy regarding the applied gain reduction. Additionally, the plot employs a linear amplitude scale; therefore, the mismatches’ perceptibility does not directly correlate with their visual representation in the plots. The LSTM model does not compress the signal promptly in the LA-2A scenario, while the other model better emulates the timing behavior of the compression process, particularly evident when observing peak reduction at $40$. Meanwhile, the S6 model significantly outperforms other models with the mildest compression applied. Conversely, when looking at the CL 1B case, although the S6 model has better matching, all the models overcompress the signal, especially the models using LSTM and S4D in the recurrent block of the proposed architecture. The ED model does not properly reduce the dynamic but similarly presents the compression as constant.
Figure 10: Spectral flux measured at the output of the Teletronix LA-2A (top) and TubeTech CL 1B (bottom) compared with those at the output of models trained using the proposed S6 and alternative architectures. The applied compression settings are indicated above the plots. The legend associates the different color curves with their corresponding models, with ’target’ denoting the analog hardware device.
@@ 5.1 Unseen settings and sampling density
The results of the experiments to evaluate the ability of models to predict the compressor output for unseen parameter combinations confirm that the proposed S6 model performs better than alternative and baseline architectures. These experiments use the Universal Audio LA-2A, Softube CL 1B ’dynamic,’ and Softube CL 1B ’time’ datasets, assess the impact of the dataset size, with respect to the control parameter sampling density on the learning accuracy, as described in Section 4.3.1. In particular, results for each dataset are summarized in Table 5, Table 6, and Table 7, detailing the metrics obtained from datasets of different sizes due to varied parameter sampling density.
@@
The results for the LA-2A dataset confirm that the S6 model has better performance. When the dataset size increases, the S6 model accuracy increases as well, together with the gap between the other proposed models. Only the ED model presents slightly and negligible better errors for the smaller size. Noticeably, all other models resulted in higher errors compared to the smaller case. Indeed, having more examples of parameter combinations does not lead to lower errors for all the models. In particular, ED, S4D, and LSTM behaved better when trained with the fewest number of examples. Similar behavior is encountered with the Softube CL 1B dynamic dataset. The S6 model is not always ranked first, but the metrics are very close to the best ones in all the scenarios, and it achieves the best absolute accuracy with the medium-size dataset. ED again shows better predictions for unseen combinations in the smaller dataset. Still increasing, dataset S6 becomes more accurate than other models and than the previous case, which is the opposite of the other models. Also, in this case, the best accuracy is achieved with the medium-size dataset. Finally, the case of unseen attack and release times show the opposite trend. This time, the errors are generally bigger for all the models when increasing the dataset size, as well as for the S6 model, which results the best also in this case. The bigger errors with this dataset confirm a difficult case when including time variant dependencies.
@@
In general, S6 is more consistent and stable when changing the dataset size. When considering parameters influencing the dynamic, S6 improves when presenting more examples; in contrast, when variable temporal behavior is considered. As mentioned, we do not know the scale of the parameter ranges. We sampled the parameters with equal-spaced steps, considering a linear range. In the case of logarithmic scales, we have more examples covering the part close to one extreme and a few examples for the other one, leading to over-fitting. Results also indicate that LSTM models, especially ED, are better when interpolating unseen parameters when a few examples are available. This suggests they better encode the input information when having a relatively small amount of examples. Conversely, S6 and S4D are better when more examples are seen during training. Lastly, bigger datasets do not always lead to lower error; the smallest ones sometimes present more accurate results. This means that too many examples can confuse the network, while fewer points in the space help better approximate the different conditioning cases overall.
@@
Table 5: Loss and performance metrics for models trained on Universal Audio LA-2A datasets with different parameter sampling densities. The ’Set’ columns indicate the training sets used, corresponding to the suffix of the training set names detailed in Table 3. The ’EP’ columns report the last training epoch at which the models exhibited improvements. Models with the lowest loss are highlighted in bold, and the minimum value of each metric is in italics.
Set Model MSE MAE ESR RMSE SFE M-STFTE PSM EP
0.10 S6 $6.10\cdot 10^{-4}$ $\mathit{1.19\cdot 10^{-2}}$ $1.17\cdot 10^{-1}$ $1.10\cdot 10^{-2}$ $\mathit{1.02\cdot 10^{-2}}$ $1.86$ $0.95$ 122
S4D $8.04\cdot 10^{-4}$ $1.34\cdot 10^{-2}$ $1.54\cdot 10^{-1}$ $1.24\cdot 10^{-2}$ $1.33\cdot 10^{-2}$ $2.21$ $0.95$ 100
ED $\mathit{5.91\cdot 10^{-4}}$ $\mathit{1.19\cdot 10^{-2}}$ $\mathit{1.14\cdot 10^{-1}}$ $\mathit{1.09\cdot 10^{-2}}$ $1.37\cdot 10^{-2}$ $\mathit{1.26}$ $0.95$ 95
LSTM $6.41\cdot 10^{-4}$ $1.20\cdot 10^{-2}$ $1.23\cdot 10^{-1}$ $1.11\cdot 10^{-2}$ $1.38\cdot 10^{-2}$ $1.67$ $0.95$ 100
0.05 S6 $\mathit{5.90\cdot 10^{-4}}$ $\mathit{1.18\cdot 10^{-2}}$ $\mathit{1.13\cdot 10^{-1}}$ $\mathit{1.09\cdot 10^{-2}}$ $1.50\cdot 10^{-2}$ $\mathit{1.37}$ $0.95$ 100
S4D $1.09\cdot 10^{-3}$ $1.54\cdot 10^{-2}$ $2.10\cdot 10^{-1}$ $1.44\cdot 10^{-2}$ $\mathit{1.10\cdot 10^{-2}}$ $2.23$ $0.95$ 100
ED $7.42\cdot 10^{-4}$ $1.28\cdot 10^{-2}$ $1.43\cdot 10^{-1}$ $1.21\cdot 10^{-2}$ $1.47\cdot 10^{-2}$ $1.86$ $0.94$ 72
LSTM $4.76\cdot 10^{-3}$ $3.67\cdot 10^{-2}$ $9.18\cdot 10^{-1}$ $3.66\cdot 10^{-2}$ $1.46\cdot 10^{-2}$ $1.96$ $0.90$ 19
0.01 S6 $\mathit{5.70\cdot 10^{-4}}$ $\mathit{1.20\cdot 10^{-2}}$ $\mathit{1.09\cdot 10^{-1}}$ $\mathit{1.12\cdot 10^{-2}}$ $1.03\cdot 10^{-2}$ $\mathit{1.21}$ $0.94$ 105
S4D $4.68\cdot 10^{-3}$ $3.60\cdot 10^{-2}$ $9.02\cdot 10^{-1}$ $3.58\cdot 10^{-2}$ $1.08\cdot 10^{-2}$ $2.98$ $0.92$ 14
ED $4.43\cdot 10^{-3}$ $3.42\cdot 10^{-2}$ $8.34\cdot 10^{-1}$ $3.39\cdot 10^{-2}$ $\mathit{9.92\cdot 10^{-3}}$ $2.97$ $0.92$ 4
LSTM $4.77\cdot 10^{-3}$ $3.66\cdot 10^{-2}$ $9.19\cdot 10^{-1}$ $3.64\cdot 10^{-2}$ $1.25\cdot 10^{-2}$ $2.98$ $0.90$ 15
@@
Table 6: Loss and performance metrics for models trained on Softube CL 1B dynamic datasets with different parameter sampling densities. The ’Set’ columns indicate the training sets used, corresponding to the suffix of the training set names detailed in Table 3. The ’EP’ columns report the last training epoch at which the models exhibited improvements. Models with the lowest loss are highlighted in bold, and the minimum value of each metric is in italics.
Set Model MSE MAE ESR RMSE SFE M-STFTE PSM EP
0.20 S6 $1.05\cdot 10^{-4}$ $6.80\cdot 10^{-3}$ $2.12\cdot 10^{-1}$ $6.50\cdot 10^{-3}$ $\mathit{1.02\cdot 10^{-2}}$ $1.86$ $0.94$ 36
S4D $1.60\cdot 10^{-4}$ $8.41\cdot 10^{-3}$ $3.25\cdot 10^{-1}$ $8.07\cdot 10^{-3}$ $1.33\cdot 10^{-2}$ $2.21$ $0.93$ 28
ED $\mathit{6.99\cdot 10^{-5}}$ $\mathit{5.55\cdot 10^{-3}}$ $\mathit{1.41\cdot 10^{-1}}$ $\mathit{5.22\cdot 10^{-3}}$ $1.37\cdot 10^{-2}$ $1.26$ $0.92$ 60
LSTM $9.29\cdot 10^{-5}$ $6.27\cdot 10^{-3}$ $1.87\cdot 10^{-1}$ $5.97\cdot 10^{-3}$ $1.38\cdot 10^{-2}$ $1.67$ $0.91$ 46
0.10 S6 $\mathit{5.41\cdot 10^{-5}}$ $\mathit{4.81\cdot 10^{-3}}$ $\mathit{1.09\cdot 10^{-1}}$ $\mathit{4.56\cdot 10^{-3}}$ $1.30\cdot 10^{-2}$ $\mathit{1.14}$ $0.92$ 36
S4D $8.74\cdot 10^{-5}$ $6.27\cdot 10^{-3}$ $1.76\cdot 10^{-1}$ $5.91\cdot 10^{-3}$ $\mathit{9.69\cdot 10^{-3}}$ $1.51$ $0.96$ 4
ED $7.60\cdot 10^{-5}$ $5.73\cdot 10^{-3}$ $1.53\cdot 10^{-1}$ $5.47\cdot 10^{-3}$ $1.27\cdot 10^{-2}$ $1.56$ $0.94$ 77
LSTM $1.75\cdot 10^{-4}$ $9.06\cdot 10^{-3}$ $3.53\cdot 10^{-1}$ $8.87\cdot 10^{-3}$ $1.62\cdot 10^{-2}$ $2.36$ $0.90$ 83
0.05 S6 $\mathit{6.46\cdot 10^{-5}}$ $\mathit{5.28\cdot 10^{-3}}$ $\mathit{1.30\cdot 10^{-1}}$ $\mathit{5.03\cdot 10^{-3}}$ $1.50\cdot 10^{-2}$ $\mathit{1.37}$ $0.92$ 73
S4D $1.57\cdot 10^{-4}$ $8.46\cdot 10^{-3}$ $3.17\cdot 10^{-1}$ $8.18\cdot 10^{-3}$ $\mathit{1.10\cdot 10^{-2}}$ $2.23$ $0.94$ 34
ED $1.02\cdot 10^{-4}$ $6.77\cdot 10^{-3}$ $2.08\cdot 10^{-1}$ $6.58\cdot 10^{-3}$ $1.47\cdot 10^{-2}$ $1.86$ $0.92$ 55
LSTM $1.12\cdot 10^{-4}$ $7.16\cdot 10^{-3}$ $2.27\cdot 10^{-1}$ $6.87\cdot 10^{-3}$ $1.46\cdot 10^{-2}$ $1.96$ $0.92$ 66

### v3

@@ 4.3.2 Listening Tests
We use Mushra tests (BS,, 2015) to evaluate the perceptual similarity of the compressed signal produced by different models against the actual reference recording taken from the analog optical compressors. Listening tests are only carried out for models trained on the TubeTech CL 1B and Teletronix LA-2A datasets. The tests are conducted using the WebMUSHRA platform  (Schoeffler et al.,, 2018) and involve only participants with music performance or production expertise. For the test, we utilize a standardized setup consisting of a laptop, studio-grade audio interface, and high-end studio headphones.
Initially, the participants are briefed on the test’s objectives and given instructions on using the listening test’s graphical user interface. Each session comprised six different listening tests conducted in a fixed order. Participants are asked to rate the following eight signals on a scale from 00 to $100$ in comparison to the reference recording: the original ‘hidden’ reference signal, the reference signal with low-pass filtering at $7$ kHz and $3.5$ kHz, the outputs of the proposed model with the S6 block, alternative models with recurrent blocks based on S4D, ED, and LSTM, and the output of commercial software emulation of the specific compressor, denoted as SW. The test signals, lasting 5 seconds each, covered a variety of drums, guitar, and piano sounds. Participants are allowed to listen to the audio samples repeatedly to ensure accurate evaluation.
Of the six tests, three pertained to models of the Teletronix LA-2A compressor and three to the TubeTech CL 1B. Baseline models were excluded after pilot listening tests as their ratings were significantly lower than those of other models, a finding expected by informal listening and the higher loss values listed in Table 5. The initial inclusion of baseline models also led to longer test durations and detrimental listening fatigue among participants since the number of stimuli approached the recommended maximum  (BS,, 2015).
@@ 5 RESULTS AND ANALYSIS
Table 5 reports the results for models trained on the Teletronix LA-2A and TubeTech CL 1B datasets, detailing loss and quantitative evaluation metrics based on test set performance. The table clearly shows that the proposed model with S6 layers outperforms all others in replicating analog optical compressors’ time-variant nonlinear sound alteration processes—processes that vary according to each device’s specific control parameters. This is particularly true for the TubeTech CL 1B, where the performance gap between the S6 and all the other models is notably larger.
In contrast, for the Teletronix LA-2A, the S6 model performs more closely to the S4D, ED, and LSTM models. In both cases, the proposed S6 model achieves lower errors for all considered metrics, with the exception of PSM, which is less indicative due to values close to $1$ for all the models. This suggests minimal differences among the models according to the PEMO-Q auditory perception model and underscores the need for actual listening tests. While S4D and ED models exhibit similar behavior on this modeling task, the LSTM models produced the worst results. A similar trend is observed with LSTM-b among the baselines. Lastly, ED-b demonstrates notably high accuracy for the Teletronix LA-2A dataset.
@@
For the MUSHRA listening test, a total of $20$ individuals participated, with no participant meeting the criteria for result exclusion. The results are illustrated in Figure 6, where the box plots display the median, lower and upper quartiles, maximum and minimum values, and outliers of the responses. In the plots, the ratings for anchors have not been included to facilitate visual comparison of the models being evaluated. Results indicate that S6 models are slightly more perceptual accurate than others, offering the closest match to the reference. This is demonstrated by the median level and the narrower interquartile range, suggesting consistent performance across listening tests. The S4D models appear to be the second best, while the LSTM models performed particularly poorly when modeling the TubeTech CL 1B, likely due to noticeable audible artifacts.
The ratings for compressed signals from the Universal Audio LA-2A and Softube CL 1B software emulations (SW), which are slightly lower than those of most models, require additional considerations. On one hand, the compression characteristics of these commercial products—details of which remain undisclosed—might differ from those produced by their actual analog counterparts, yet they are still considered sufficiently good for sound and music production. During the MUSHRA tests, participants were asked to rate the similarity to a reference signal. The complete absence of background noise in digital plugins, which is typically present in small amounts in recordings from analog devices, may have also contributed to their lower similarity ratings. Further studies are needed to confirm the validity and determine the respective weight of these hypotheses.
@@
For further qualitative evaluation, we have computed a collection of plots to compare the response of the Teletronix LA-2A and TubeTech CL 1B, used as targets for the trainings, with that of the proposed S6 model and alternative models with recurrent blocks based on S4D, ED, and LSTM. These plots were generated for a selection of compression parameter settings representing cases commonly used in real application scenarios. Due to space limitations, this paper only includes those related to the following settings: Teletronix LA-2A in limit mode with peak reduction set to $20$ and compression mode with peak reduction set to $40$; TubeTech CL 1B with a threshold $-20$ dBu and ratio $2:1$, and with threshold $-10$ dBu and ratio $6:1$, both with attack and release set at $\sim 250$ ms and $\sim 5$ s.
The plots we computed include the spectrograms of the output signals in Figure 7 and Figure 8, the RMS of the output providing an indication of the signal level over time in Figure 9, and the spectral flux of the output highlighting the response during sharp attacks in Figure 10. Spectrograms are calculated using windows of $1024$ samples with $25\%$ overlap. The RMS and spectral flux are computed using windows of $4096$ samples with $25\%$ overlap. Additional plots for more compression cases and the associated target and model-generated audio signals are available online  [^6] https://github.com/RiccardoVib/Optical-DRC-with-Selective-SSMs.
Figure 7: Spectrograms of the output of the Teletronix LA-2A and of the output of models, trained using the proposed S6 and alternative architectures. Each column presents spectrograms for different compression settings labeled above the columns. The model that produced each spectrogram is specified at the top of the respective spectrogram. ‘Target’ refers to the analog hardware device.
Figure 8: Spectrograms of the output of the TubeTech CL 1B, and of the output of models trained using the proposed S6 and alternative architectures. Each column presents spectrograms for different compression settings labeled above the columns. The model that produced each spectrogram is specified at the top of the respective spectrogram. ‘Target’ refers to the analog hardware device.
@@
Figure 9: RMS levels measured at the output of the Teletronix LA-2A (top) and TubeTech CL 1B (bottom) compared with those at the output of models trained using the proposed S6 and alternative architectures. The applied compression settings are indicated above the plots. The legend associates the different color curves with their corresponding models, with ‘Target’ denoting the analog hardware device.
In replicating the appropriate gain reduction required to match the target compressed signal, the LSTM also shows the poorest performance, as evident in the RMSE plot examples of Figure 9, which refers to a $3$-second snippet from the test set. The snippet features two plucked electric bass notes in the first two seconds, followed by vocals in the final second. The first bass note is muted, while the second, occurring in the middle of the snippet, has a longer sustain. These examples have been selected as they clearly illustrate the differing behaviors of various models compared to the target, especially during the sustained bass note, which presents the greatest challenge, representing a region of the test set with errors notably higher than the average. The concluding vocal segment displays a series of rapid, albeit less pronounced, level changes that several models fail to compress accurately.
It appears that most models follow the amplitude envelope of the output signal, yet they vary in their degree of accuracy regarding the applied gain reduction. Additionally, the plot employs a linear amplitude scale; therefore, the mismatches’ perceptibility does not directly correlate with their visual representation in the plots. The LSTM model does not compress the signal promptly in the LA-2A scenario, while the other model better emulates the timing behavior of the compression process, particularly evident when observing peak reduction at $40$. Meanwhile, the S6 model significantly outperforms other models with the mildest compression applied. Conversely, when looking at the CL 1B case, although the S6 model has better matching, all the models overcompress the signal, especially the models using LSTM and S4D in the recurrent block of the proposed architecture. The ED model does not properly reduce the dynamic but similarly presents the compression as constant.
Figure 10: Spectral flux measured at the output of the Teletronix LA-2A (top) and TubeTech CL 1B (bottom) compared with those at the output of models trained using the proposed S6 and alternative architectures. The applied compression settings are indicated above the plots. The legend associates the different color curves with their corresponding models, with ‘Target’ denoting the analog hardware device.
@@ 5.1 Unseen settings and sampling density
The results of the experiments to evaluate the ability of models to predict the compressor output for unseen parameter combinations confirm that the proposed S6 model performs better than alternative and baseline architectures. These experiments use the Universal Audio LA-2A, Softube CL 1B ‘dynamic,’ and Softube CL 1B ‘time’ datasets, assess the impact of the dataset size, with respect to the control parameter sampling density on the learning accuracy, as described in Section 4.3.1. The performance of the models is always evaluated on the same test set; results for each dataset are summarized in Table 6, Table 7, and Table 8, which detail the metrics across datasets of varying sizes due to differences in parameter sampling density.
@@
The results for the LA-2A dataset confirm that the S6 model has better performance. When the dataset size increases, the S6 model accuracy increases as well, together with the gap between the other proposed models. Only the ED model presents slightly and negligible better errors for the smaller size. Noticeably, all other models resulted in higher errors compared to the smaller case. Indeed, having more examples of parameter combinations does not lead to lower errors for all the models. In particular, ED, S4D, and LSTM behaved better when trained with the fewest number of examples. Similar behavior is encountered with the Softube CL 1B dynamic dataset.
The S6 model is not always ranked first, but the metrics are very close to the best ones in all the scenarios, and it achieves the best absolute accuracy with the medium-size dataset. ED again shows better predictions for unseen combinations in the smaller dataset. Still increasing, dataset S6 becomes more accurate than other models and than the previous case, which is the opposite of the other models. Also, in this case, the best accuracy is achieved with the medium-size dataset.
Finally, the case of unseen attack and release times show the opposite trend. This time, the errors are generally larger for all models when increasing the dataset size, as well as for the S6 model, which results in the best performing one also in this case. The bigger errors with this dataset confirm a difficult case when including time variant dependencies.
@@
In general, S6 is more consistent and stable when changing the dataset size. When considering parameters that influence the dynamics, S6’s performance improves with an increased number of input-output examples; in contrast, a decline is observed when dealing with data characterized by variable temporal behavior. As mentioned, we do not know the scale of the parameter ranges. We sampled the parameters with equal-spaced steps, considering a linear range. In the case of logarithmic scales, we have more examples covering the part close to one extreme and a few examples for the other one, leading to over-fitting. Results also indicate that LSTM models, especially ED, are better when interpolating unseen parameters when a few examples are available. This suggests they better encode the input information when having a relatively small amount of examples. Conversely, S6 and S4D are better when more examples are seen during training.
Lastly, bigger datasets do not always lead to lower error; the smallest ones sometimes present more accurate results. This means that too many examples can confuse the network, while fewer points in the space help better approximate the different conditioning cases overall.
@@
Table 6: Loss and performance metrics for models trained on Universal Audio LA-2A datasets with different parameter sampling densities. The ‘Set’ columns indicate the training sets used, corresponding to the suffix of the training set names detailed in Table 3. Models with the lowest loss and the minimum value for each metric are highlighted in bold.
Set Model MSE MAE ESR RMSE SFE M-STFTE PSM
0.10 S6 $6.10\cdot 10^{-4}$ $\mathbf{1.19\cdot 10^{-2}}$ $1.17\cdot 10^{-1}$ $1.10\cdot 10^{-2}$ $\mathbf{1.02\cdot 10^{-2}}$ $1.86$ $0.95$
S4D $8.04\cdot 10^{-4}$ $1.34\cdot 10^{-2}$ $1.54\cdot 10^{-1}$ $1.24\cdot 10^{-2}$ $1.33\cdot 10^{-2}$ $2.21$ $0.95$
ED $\mathbf{5.91\cdot 10^{-4}}$ $\mathbf{1.19\cdot 10^{-2}}$ $\mathbf{1.14\cdot 10^{-1}}$ $\mathbf{1.09\cdot 10^{-2}}$ $1.37\cdot 10^{-2}$ $\mathbf{1.26}$ $0.95$
LSTM $6.41\cdot 10^{-4}$ $1.20\cdot 10^{-2}$ $1.23\cdot 10^{-1}$ $1.11\cdot 10^{-2}$ $1.38\cdot 10^{-2}$ $1.67$ $0.95$
0.05 S6 $\mathbf{5.90\cdot 10^{-4}}$ $\mathbf{1.18\cdot 10^{-2}}$ $\mathbf{1.13\cdot 10^{-1}}$ $\mathbf{1.09\cdot 10^{-2}}$ $1.50\cdot 10^{-2}$ $\mathbf{1.37}$ $0.95$
S4D $1.09\cdot 10^{-3}$ $1.54\cdot 10^{-2}$ $2.10\cdot 10^{-1}$ $1.44\cdot 10^{-2}$ $\mathbf{1.10\cdot 10^{-2}}$ $2.23$ $0.95$
ED $7.42\cdot 10^{-4}$ $1.28\cdot 10^{-2}$ $1.43\cdot 10^{-1}$ $1.21\cdot 10^{-2}$ $1.47\cdot 10^{-2}$ $1.86$ $0.94$
LSTM $4.76\cdot 10^{-3}$ $3.67\cdot 10^{-2}$ $9.18\cdot 10^{-1}$ $3.66\cdot 10^{-2}$ $1.46\cdot 10^{-2}$ $1.96$ $0.90$
0.01 S6 $\mathbf{5.70\cdot 10^{-4}}$ $\mathbf{1.20\cdot 10^{-2}}$ $\mathbf{1.09\cdot 10^{-1}}$ $\mathbf{1.12\cdot 10^{-2}}$ $1.03\cdot 10^{-2}$ $\mathbf{1.21}$ $0.94$
S4D $4.68\cdot 10^{-3}$ $3.60\cdot 10^{-2}$ $9.02\cdot 10^{-1}$ $3.58\cdot 10^{-2}$ $1.08\cdot 10^{-2}$ $2.98$ $0.92$
ED $4.43\cdot 10^{-3}$ $3.42\cdot 10^{-2}$ $8.34\cdot 10^{-1}$ $3.39\cdot 10^{-2}$ $\mathbf{9.92\cdot 10^{-3}}$ $2.97$ $0.92$
LSTM $4.77\cdot 10^{-3}$ $3.66\cdot 10^{-2}$ $9.19\cdot 10^{-1}$ $3.64\cdot 10^{-2}$ $1.25\cdot 10^{-2}$ $2.98$ $0.90$
@@
Table 7: Loss and performance metrics for models trained on Softube CL 1B dynamic datasets with different parameter sampling densities. The ‘Set’ columns indicate the training sets used, corresponding to the suffix of the training set names detailed in Table 3. Models with the lowest loss and the minimum value for each metric are highlighted in bold.
Set Model MSE MAE ESR RMSE SFE M-STFTE PSM
0.20 S6 $1.05\cdot 10^{-4}$ $6.80\cdot 10^{-3}$ $2.12\cdot 10^{-1}$ $6.50\cdot 10^{-3}$ $\mathbf{1.02\cdot 10^{-2}}$ $1.86$ $0.94$
S4D $1.60\cdot 10^{-4}$ $8.41\cdot 10^{-3}$ $3.25\cdot 10^{-1}$ $8.07\cdot 10^{-3}$ $1.33\cdot 10^{-2}$ $2.21$ $0.93$
ED $\mathbf{6.99\cdot 10^{-5}}$ $\mathbf{5.55\cdot 10^{-3}}$ $\mathbf{1.41\cdot 10^{-1}}$ $\mathbf{5.22\cdot 10^{-3}}$ $1.37\cdot 10^{-2}$ $\mathbf{1.26}$ $0.92$
LSTM $9.29\cdot 10^{-5}$ $6.27\cdot 10^{-3}$ $1.87\cdot 10^{-1}$ $5.97\cdot 10^{-3}$ $1.38\cdot 10^{-2}$ $1.67$ $0.91$
0.10 S6 $\mathbf{5.41\cdot 10^{-5}}$ $\mathbf{4.81\cdot 10^{-3}}$ $\mathbf{1.09\cdot 10^{-1}}$ $\mathbf{4.56\cdot 10^{-3}}$ $1.30\cdot 10^{-2}$ $\mathbf{1.14}$ $0.92$
S4D $8.74\cdot 10^{-5}$ $6.27\cdot 10^{-3}$ $1.76\cdot 10^{-1}$ $5.91\cdot 10^{-3}$ $\mathbf{9.69\cdot 10^{-3}}$ $1.51$ $0.96$
ED $7.60\cdot 10^{-5}$ $5.73\cdot 10^{-3}$ $1.53\cdot 10^{-1}$ $5.47\cdot 10^{-3}$ $1.27\cdot 10^{-2}$ $1.56$ $0.94$
LSTM $1.75\cdot 10^{-4}$ $9.06\cdot 10^{-3}$ $3.53\cdot 10^{-1}$ $8.87\cdot 10^{-3}$ $1.62\cdot 10^{-2}$ $2.36$ $0.90$
0.05 S6 $\mathbf{6.46\cdot 10^{-5}}$ $\mathbf{5.28\cdot 10^{-3}}$ $\mathbf{1.30\cdot 10^{-1}}$ $\mathbf{5.03\cdot 10^{-3}}$ $1.50\cdot 10^{-2}$ $\mathbf{1.37}$ $0.92$
S4D $1.57\cdot 10^{-4}$ $8.46\cdot 10^{-3}$ $3.17\cdot 10^{-1}$ $8.18\cdot 10^{-3}$ $\mathbf{1.10\cdot 10^{-2}}$ $2.23$ $0.94$
ED $1.02\cdot 10^{-4}$ $6.77\cdot 10^{-3}$ $2.08\cdot 10^{-1}$ $6.58\cdot 10^{-3}$ $1.47\cdot 10^{-2}$ $1.86$ $0.92$
LSTM $1.12\cdot 10^{-4}$ $7.16\cdot 10^{-3}$ $2.27\cdot 10^{-1}$ $6.87\cdot 10^{-3}$ $1.46\cdot 10^{-2}$ $1.96$ $0.92$